OpenStudy (anonymous):

Find the integral of dx/x^(1.01) from 1 to infinity.

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

\(\Large x^{1.01}=(x^1)^{\frac{1}{100}}x=x^{\frac{101}{100}} = \sqrt[100]{x}^{101} \)

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty}x ^{1.01}dx=\frac{ x ^{2.01} }{ 2.01 }~from~1~\to~\infty approaches~\to~\infty \]\]because x>1

OpenStudy (anonymous):

Isn't it 1/x^(1.01) dx?

OpenStudy (anonymous):

no matter how you look at it, both answers demonstrate that you're integrating a polynomial, and you know how to integrate these I presume. Use the regular formula.

OpenStudy (anonymous):

i forgot to see / sign in that case write -1.01 in the exponent

OpenStudy (anonymous):

Never mind, I got it by using the regular formula. Thank you, Spacelimbus.

OpenStudy (anonymous):

after solving answer no infinity some real no.you can solve it ,if any problem i will help you.